OpenStudy (anonymous):

Write the equation of a circle whose center is at the origin and that contains the point (-3, 0).

OpenStudy (ryan123345):

ok first you have to use the distance formula to find the radius.

OpenStudy (nurali):

x^2+y^2=r^2 x^2+y^2=9